# Van der Waals devices for surface-sensitive experiments

### v1

@@
Abstract
@@
In-operando characterization of van der Waals (vdW) devices using surface-sensitive methods provides critical insights into phase transitions and correlated electronic states. Yet, integrating vdW materials in functional devices while maintaining pristine surfaces is a key challenge for combined transport and surface-sensitive experiments. Conventional lithographic techniques introduce surface contamination, limiting the applicability of state-of-the-art spectroscopic probes. We present a stencil lithography-based approach for fabricating vdW devices, producing micron-scale electrical contacts, and exfoliation in ultra-high vacuum. The resist-free patterning method utilizes a shadow mask to define electrical contacts and yields thin flakes down to the single-layer regime via gold-assisted exfoliation. As a demonstration, we fabricate devices from 1T–TaS2 flakes, achieving reliable contacts for application of electrical pulses and resistance measurements, as well as clean surfaces allowing for angle-resolved photoemission spectroscopy. The approach provides a platform for studying the electronic properties of vdW systems with surface-sensitive probes in well-defined device geometries.
I Introduction
@@
Van der Waals (vdW) materials provide a rich playground for exploring correlated electronic states and phase transitions in two-dimensional systems, with applications in next-generation electronic devices based on thin vdW flakes. They enable the exploration of novel quantum states with unprecedented control via gating and pulse application [1, 2, 3, 4, 5]. While transport measurements provide insight into the electronic properties of such devices, spectroscopic tools are necessary to directly probe the electronic band structure. Combining these approaches remains challenging, as conventional lithographic device fabrication methods introduce surface contamination or modify the surface, preventing the use of state-of-the-art surface-sensitive tools, such as scanning tunneling microscopy [6] or angle-resolved photoemission spectroscopy (ARPES) [7]. Thus, a critical challenge in fabricating devices for surface-sensitive experiments involving vdW materials is achieving three aspects simultaneously: (i) obtaining thin, uniform flakes down to monolayers, (ii) integrating electrical contacts at the micron-scale for well-defined device geometries, and (iii) maintaining pristine surface quality through in-situ exfoliation in ultra-high vacuum (UHV).
@@
with pre-attached contacts [12, 13] |
| Glovebox device assembly [14, 15, 16, 17] | ✓ | ✓ | ✗∗ |
| UHV fabrication system [18] | ✓ | ✗ | ✓ |
| Gold-assisted exfoliation [19, 20, 21, 22] | ✓ | ✗† | ✓ |
@@
A variety of fabrication methods have been developed to integrate vdW materials into devices, but most approaches satisfy only two out of the three requirements above (see Tab. 1). Fabrication methods involving standard dry pick-up transfer techniques with viscoelastic polymer stamps, e.g. polydimethylsiloxane (PDMS) [8], are widely used for assembling vdW heterostructures, but they often leave polymer residues and water on the surface, which degrade the interface quality and limit compatibility with surface-sensitive techniques [23, 24]. While such contamination can in principle be removed, e.g. from graphene devices by annealing at C [9], this approach is viable only for certain “robust” vdW materials. In many cases, such treatment leads to irreversible structural changes. For instance, for the prototypical transition metal dichalcogenide material 1T–TaS2, annealing at C induces a polytype transition to the semimetallic 2H phase [25]. Alternative strategies to avoid polymer contamination include protective capping with hBN or graphene [10] during transfer or the use of metalized SiNx membranes [11]. Though, both approaches still require mild annealing and are difficult to integrate with patterned device architectures.
Cleaving bulk vdW crystals in UHV with pre-attached contacts has recently also been used in transport-ARPES studies on Ca2RuO4 [12] and 1T–TaS2 [13]. This method exposes a pristine surface in-situ but is limited to bulk materials, making it incompatible with micron-scale device designs.
Fabrication in inert atmospheres such as in N2 or Ar gloveboxes [14, 15, 16, 17], or even under UHV [18], has been explored to exfoliate thin flakes and assemble heterostructures while minimizing oxidation. However, even within a controlled N2 environment, water adsorption on the surface remains an issue, often necessitating an annealing step to obtain clean surfaces. Moreover, such fabrication setups require specialized equipment that is rarely available near UHV measurement facilities.
Finally, gold-assisted exfoliation [19, 20, 21, 22] allows for preparation of large-area monolayer flakes with clean surfaces by leveraging the strong interaction between freshly evaporated Au and chalcogen atoms. While Au is most commonly used due to its inert nature, it can be replaced by other metals such as Pd, Ni, Cu, and Ag, which exhibit similarly high adhesion through strong bonding to chalcogen atoms [26]. Importantly, the timing of exfoliation is critical: contaminants adsorb on Au within minutes, making the surface hydrophobic, which significantly reduces the monolayer yield [27, 21]. This degradation cannot be reversed by cleaning or annealing, precluding post-processing steps such as metal lift-off. While the method yields high-quality flakes, device integration typically involves patterning steps after exfoliation, for example, by etching in potassium iodide solution [22].
@@
As a demonstration of this technique, we fabricate two-terminal devices using 1T–TaS2, a prototypical vdW material due to its rich electronic phase diagram, including a Mott insulating low-temperature state [28, 29, 30, 31, 32], different charge-density wave (CDW) phases [33, 34], superconductivity upon pressure [35], and a putative quantum spin liquid state [36, 37]. Furthermore, 1T–TaS2 exhibits a non-equilibrium metallic “hidden” state that can be reversibly induced by optical or electrical pulses [4, 38, 39, 40, 41]. This nonvolatile state holds promise for use in the next generation of cryo-memory devices due to energy efficiency and switching speed [42, 43, 44].
@@
The devices presented here are tailored for micro-beam ARPES [45, 46] in combination with transport measurements and application of electrical (switching) pulses. More broadly, the fabrication method enables studies of the electronic properties of thin vdW flakes and provides a platform for integrating surface-sensitive techniques with transport measurements across a wide range of material systems.
II Results and discussion
II.1 Device fabrication
@@
Here, we outline the steps for the preparation of the vdW device. Details of the shadow-mask fabrication are provided in the Methods section. To benefit from the high adhesion of gold-assisted exfoliation, it is crucial to fabricate electrodes in a way that avoids contamination. This can be achieved through evaporation of Au through a shadow mask [47, 48, 49, 50]. Such apertures can be made using various techniques, such as mechanical cutting, focused ion beam milling [51, 52, 53], or etching into Si wafers [54, 55, 56], with the choice depending on the required feature size. For our approach, we employ laser lithography combined with deep reactive ion etching (DRIE) [57] to produce high-aspect ratio masks with well-defined micron-scale features (see Fig. 2a and the Methods). This stencil wafer is placed on the Si/SiO2 substrate during the deposition of Ti/Au contacts via electron beam evaporation. After the contact deposition, bulk 1T–TaS2 crystals with typical dimensions of mm in diameter are cleaved on tape loops, positioned over the gap between the Au pads, brought into contact with the substrate, and gently pressed to ensure adhesion.
@@
Due to rapid degradation of the Au surface in air, it is crucial that this is performed within min after venting the evaporation chamber to obtain a high yield of thin flakes [21]. Next, electrical connections are made by attaching wires to the contact pads using Ag epoxy. A fully assembled device is shown in Fig. S1.
@@
To obtain a contamination-free surface, the exfoliation process is performed later under UHV conditions by removing the tape loop with a mechanical manipulator. This process cleaves the vdW crystal, leaving flakes with freshly exposed surfaces on the contacts at random positions; only flakes that bridge the gap between the contact pads form functional devices. Electrical connectivity is tested and the flakes are visually inspected using an optical microscope, as well as in our ARPES demonstration the spectral intensity of the Ta 4 core levels, mainly to confirm flake dimensions, thickness, and surface homogeneity (see Fig. 2b). The vdW flakes, like the one shown in Fig. 2c, have lateral dimensions from tens to hundreds of microns and thicknesses ranging from a single monolayer up to 100 nm, enabling access to both few-layer and bulk-like properties.
II.2 Device characterization
@@
Figure 2d shows the temperature dependence of the two-terminal resistance upon cooling after exfoliation at 250 K, including the metal-insulator transition from the the nearly-commensurate (NC) to commensurate (C)CDW phase of 1T–TaS2. Large contact resistances on the order of several hundreds of are typically observed when vdW materials are probed with metal contacts on their flat face [24]. However, we observe exceptionally low contact resistances on the order of in all our devices, e.g. the two-terminal resistance of the device shown in Fig. 2c is at room temperature, which we attribute to the large contact area with covalent-like bonding at the 1T–TaS2 / Au interface [22]. The high-quality contacts allow for low-voltage state switching of the 1T–TaS2 device, where electrical excitation induces a phase transition from the insulating, equilibrium CCDW phase to a metallic, nonthermal and metastable “hidden” state [38]. This phase transition is characterized by a drop in resistance and associated with the disruption of long-range CCDW order. Recent studies suggest that the emergence of the “hidden” state is driven by out-of-plane restacking and the disappearance of interlayer dimerization of the CCDW phase [39, 40, 41, 32]. In our case, at K, we observe a resistance drop from 186 to upon application of a V, µs voltage pulse (see Fig. 2e). Although the switching to the “hidden” state is apparent from the transport measurements, the short lifetime of this metastable phase at K prevents direct detection by techniques with longer acquisition times, such as ARPES [42].
@@
Angle-resolved photoemission spectroscopy: Figure 3 summarizes spatially-resolved ARPES measurements performed on the device shown in Fig. 2, using a photon energy of eV and a beam spot focused to µm2. Figure 3b shows representative angle-integrated and momentum-resolved spectra of this flake. In particular, the dispersion of the main valence bands, formed by S orbitals [34], is well resolved (see also Fig. S2). This is further validated by a constant energy contour at eV with details and resolution comparable to those in studies on bulk crystals [58, 59, 60, 40, 61].
@@
A spatially resolved Ta core-level scan with a step size of µm is used to group the data points based on spectral similarity by the -means clustering algorithm [62] (see Fig. 3d). This unsupervised learning algorithm partitions the dataset into clusters by minimizing the intra-cluster variance. We determine the optimal using the “elbow” method, where diminishing returns in reducing the within-cluster sum of squares are observed. Here, we obtain (see Fig. S3), where the three clusters correspond to distinct regions of the scanned area: bulk ( layers) and monolayer 1T–TaS2 flake, as well as Au surface, agreeing well with the optical microscope image (see Fig. 3a and 3d, as well as the overlaid image in Fig. S4). All angle-integrated spectra within a cluster are averaged to obtain the mean and standard deviation shown in Fig. 3e. A Shirley-type background [63] is subtracted to account for inelastic scattering (see Fig. S5).
@@
Within the bulk-like cluster, the Ta core levels exhibit a splitting stemming from the different electronic environments of Ta atoms within the star-of-David-shaped lattice distortion in the CDW phase [33]. As shown in the inset in Fig. 3e, there are three distinct Ta sites, namely a central (A), as well as an inner (B) and outer (C) ring position. Due to their different electronic environment and the charge redistribution towards the central A site, they appear at different binding energies, i.e. eV (C), eV (B), and eV (A) in accordance with measurements of bulk single-crystals [64]. The ratio of the intensities matches that of the Ta atoms of each species, i.e. ABC 166. We attribute deviations of the fit from the experimental data, most pronounced for the C species, to inequivalent C sites due to the complex out-of-plane stacking [31, 32], or puckering distortions [65] present in the CDW states of 1T–TaS2. The discrepancy at the edge of the spectrum can be attributed to systematic errors from the background subtraction procedure.
@@
In contrast, the spectrum taken from the monolayer cluster shows a reduction of the splitting between the B and C species, alongside a reduced relative C intensity, which we attribute to the absence of out-of-plane stacking. A shift of eV towards lower binding energies is observed, which we ascribe to charge transfer from Au to 1T–TaS2 and is consistent with findings in monolayer MoS2 on Au contacts [21]. To study the influence of 1T–TaS2 / Au interface in more detail, one could aim for a device with a 1T–TaS2 monolayer that extends over the gapped region over the substrate in a future experiment, allowing for a direct comparison of the Au-bound and freestanding monolayer.
III Conclusions

### v2

# ![[Uncaptioned image]](x3.png) Van der Waals devices for surface-sensitive experiments | | Nicolai Taufertshöfer,a,b Corinna Burri,a,c Rok Venturini,a Iason Giannopoulos,a Sandy Adhitia Ekahana,a Enrico Della Valle,a,c Anže Mraz,d,e Yevhenii Vaskivskyi,d,e Jan Lipic,d Alexei Barinov,f Dimitrios Kazazis,a Yasin Ekinci,a Dragan Mihailovic,∗d,e,g and Simon Gerber∗a | |

@@
![[Uncaptioned image]](x3.png)
Van der Waals devices for surface-sensitive experiments |
| Nicolai Taufertshöfer,a,b Corinna Burri,a,c Rok Venturini,a Iason Giannopoulos,a Sandy Adhitia Ekahana,a Enrico Della Valle,a,c Anže Mraz,d,e Yevhenii Vaskivskyi,d,e Jan Lipic,d Alexei Barinov,f Dimitrios Kazazis,a Yasin Ekinci,a Dragan Mihailovic,∗d,e,g and Simon Gerber∗a | |
@@
In-operando characterization of van der Waals (vdW) devices using surface-sensitive methods provides critical insights into phase transitions and correlated electronic states. Yet, integrating vdW materials in functional devices while maintaining pristine surfaces is a key challenge for combined transport and surface-sensitive experiments. Conventional lithographic techniques introduce surface contamination, limiting the applicability of state-of-the-art spectroscopic probes. We present a stencil lithography-based approach for fabricating vdW devices, producing micron-scale electrical contacts, and exfoliation in ultra-high vacuum. The resist-free patterning method utilizes a shadow mask to define electrical contacts and yields thin flakes down to the single-layer regime via gold-assisted exfoliation. As a demonstration, we fabricate devices from 1T–TaS2 flakes, achieving reliable contacts for application of electrical pulses and resistance measurements, as well as clean surfaces allowing for angle-resolved photoemission spectroscopy. The approach provides a platform for studying the electronic properties of vdW systems with surface-sensitive probes in well-defined device geometries. |
1 Introduction
@@
Van der Waals (vdW) materials provide a rich playground for exploring correlated electronic states and phase transitions in two-dimensional systems, with applications in next-generation electronic devices based on thin vdW flakes. They enable the exploration of novel quantum states with unprecedented control via gating and pulse application 1, 2, 3, 4, 5. While transport measurements provide insight into the electronic properties of such devices, spectroscopic tools are necessary to directly probe the electronic band structure. Combining these approaches remains challenging as conventional lithographic device fabrication methods introduce surface contamination or modify the surface, preventing the use of state-of-the-art surface-sensitive tools, such as scanning tunneling microscopy 6 or angle-resolved photoemission spectroscopy (ARPES) 7. Thus, a critical challenge in fabricating devices for surface-sensitive experiments involving vdW materials is achieving three aspects simultaneously: (i) obtaining thin, uniform flakes down to monolayers, (ii) integrating electrical contacts at the micron-scale for well-defined device geometries, and (iii) maintaining pristine surface quality through in-situ exfoliation in ultra-high vacuum (UHV).
A variety of fabrication methods have been developed to integrate vdW materials into devices, but most approaches satisfy only two out of the three requirements above (see Tab. 1). Fabrication methods involving standard dry pick-up transfer techniques with viscoelastic polymer stamps, e.g. polydimethylsiloxane (PDMS) 8, are widely used for assembling vdW heterostructures, but they often leave polymer residues and water on the surface, which degrade the interface quality and limit compatibility with surface-sensitive techniques 9, 10. While such contamination can in principle be removed, e.g. from graphene devices by annealing at C 11, this approach is viable only for certain “robust” vdW materials. In many cases, such treatment leads to irreversible structural changes. For instance, for the prototypical transition metal dichalcogenide material 1T–TaS2, annealing at C induces a polytype transition to the semimetallic 2H phase 12. Alternative strategies to avoid polymer contamination include protective capping with hBN or graphene 13 during transfer or the use of metalized SiNx membranes 14. Though, both approaches still require mild annealing and are difficult to integrate with patterned device architectures.
Cleaving bulk vdW crystals in UHV with pre-attached contacts has recently also been used in transport-ARPES studies on Ca2RuO4 15 and 1T–TaS2 16. This method exposes a pristine surface in-situ but is limited to bulk materials, making it incompatible with micron-scale device designs.
@@
with pre-attached contacts 15, 16 |
| Glovebox device assembly 17, 18, 19, 20 | ✓ | ✓ | ✗1 |
| UHV fabrication system 21 | ✓ | ✗ | ✓ |
| Gold-assisted exfoliation 22, 23, 24, 25 | ✓ | ✗2 | ✓ |
@@
Fabrication in inert atmospheres such as in N2 or Ar gloveboxes 17, 18, 19, 20, or even under UHV 21, has been explored to exfoliate thin flakes and assemble heterostructures while minimizing oxidation. However, even within a controlled N2 environment, water adsorption on the surface remains an issue, often necessitating an annealing step to obtain clean surfaces. Moreover, such fabrication setups require specialized equipment that is rarely available near UHV measurement facilities.
Finally, gold-assisted exfoliation 22, 23, 24, 25 allows for preparation of large-area monolayer flakes with clean surfaces by leveraging the strong interaction between freshly evaporated Au and chalcogen atoms. While Au is most commonly used due to its inert nature, it can be replaced by other metals such as Pd, Ni, Cu, and Ag, which exhibit similarly high adhesion through strong bonding to chalcogen atoms 26. Importantly, the timing of exfoliation is critical: contaminants adsorb on Au within minutes, making the surface hydrophobic, which significantly reduces the monolayer yield 27, 24. This degradation cannot be reversed by cleaning or annealing, precluding post-processing steps such as metal lift-off. While the method yields high-quality flakes, device integration typically involves patterning steps after exfoliation, for example, by etching in potassium iodide solution 25.
@@
As a demonstration of this technique, we fabricate two-terminal devices of 1T–TaS2, a prototypical vdW material due to its rich electronic phase diagram, including a Mott insulating low-temperature state 28, 29, 30, 31, 32, different charge-density wave (CDW) phases 33, 34, superconductivity upon pressure 35, and a putative quantum spin liquid state 36, 37. Furthermore, 1T–TaS2 exhibits a non-equilibrium metallic “hidden” state that can be reversibly induced by optical or electrical pulses 4, 38, 39, 40, 41. This non-volatile state holds promise for use in the next generation of cryo-memory devices due to energy efficiency and switching speed 42, 43, 44.
@@
The devices presented here are tailored for micro-beam ARPES 45, 46 in combination with transport measurements and application of electrical (switching) pulses. More broadly, the fabrication method enables studies of the electronic properties of thin vdW flakes and provides a platform for integrating surface-sensitive techniques with transport measurements across a wide range of material systems.
2 Results and discussion
2.1 Device fabrication
@@
Here, we outline the steps for the preparation of the vdW device. Details of the shadow-mask fabrication are provided in the Methods section. To benefit from the high adhesion of gold-assisted exfoliation, it is crucial to fabricate electrodes in a way that avoids contamination. This can be achieved through evaporation of Au through a shadow mask 47, 48, 49, 50. Such apertures can be made using various techniques, such as mechanical cutting, focused ion beam milling 51, 52, 53, or etching into Si wafers 54, 55, 56, with the choice depending on the required feature size. For our approach, we employ laser lithography combined with deep reactive ion etching (DRIE) 57 to produce high-aspect ratio masks with well-defined micron-scale features (see Fig. 2a and the Methods). This stencil wafer is placed on the Si/SiO2 substrate during the deposition of Ti/Au contacts via electron beam evaporation. After the contact deposition, bulk 1T–TaS2 crystals with typical dimensions of mm in diameter are cleaved on tape loops, positioned over the gap between the Au pads, brought into contact with the substrate, and gently pressed to ensure adhesion.
@@
Due to rapid degradation of the Au surface in air, it is crucial that this is performed within min after venting the evaporation chamber to obtain a high yield of thin flakes 24. Next, electrical connections are made by attaching wires to the contact pads using Ag epoxy. A fully assembled device is shown in Fig. S1.
@@
To obtain a contamination-free surface, the exfoliation process is performed later under UHV conditions by removing the tape loop with a mechanical manipulator. This process cleaves the vdW crystal, leaving flakes with freshly exposed surfaces on the contacts at random positions; only flakes that bridge the gap between the contact pads form functional devices. Electrical connectivity is tested and the flakes are visually inspected using an optical microscope, as well as in our ARPES demonstration the spectral intensity of the Ta core levels, mainly to confirm flake dimensions, thickness, and surface homogeneity (see Fig. 2b). The vdW flakes, like the one shown in Fig. 2c, have lateral dimensions from tens to hundreds of microns. Atomic force microscopy (AFM) reveals thicknesses ranging from a monolayer up to nm (see Fig. S2), enabling access to both few-layer and bulk properties. Such variation in thickness can occur even within a single macroscopic flake as local tearing and a non-uniform stress distribution during peeling, combined with the material’s in-plane stiffness, can result in different exfoliation planes. In our experiments, we routinely observe mono-, bi-, and few-layer flakes. While the exfoliation process is inherently non-deterministic, the yield and thickness distribution of flakes can be influenced by tuning the Au surface properties—such as roughness and cleanliness—through parameters like evaporation rate, film thickness, adhesion layer, and time elapsed after breaking vacuum. This enables the selection of flakes with the desired thickness from the exfoliated set.
2.2 Device characterization
@@
Figure 2d shows the temperature dependence of the two-terminal resistance upon cooling after exfoliation at 250 K, including the metal-insulator transition from the the nearly-commensurate (NC) to commensurate (C) CDW phase of 1T–TaS2. Large contact resistances on the order of several hundreds of are typically observed when vdW materials are probed with metal contacts on their flat face 10. However, we observe exceptionally low contact resistances on the order of in all our devices, e.g. the two-terminal resistance of the device shown in Fig. 2c is at room temperature, which we attribute to the large contact area with covalent-like bonding at the 1T–TaS2 / Au interface 25. The high-quality contacts allow for low-voltage state switching of the 1T–TaS2 device, where electrical excitation induces a phase transition from the insulating, equilibrium CCDW phase to a metallic, non-thermal and metastable “hidden” state 38. This phase transition is characterized by a drop in resistance and associated with the disruption of long-range CCDW order. Recent studies suggest that the emergence of the hidden state is driven by out-of-plane restacking and the disappearance of interlayer dimerization of the CCDW phase 39, 40, 41, 32. In our case, at K, we observe a resistance drop from 186 to upon application of a V, µs voltage pulse (see Fig. 2e). Although the switching to the hidden state is apparent from the transport measurements, the short lifetime of this metastable phase at K prevents direct detection by techniques with longer acquisition times, such as ARPES 42.
Angle-resolved photoemission spectroscopy:
@@
Figure 3 summarizes spatially-resolved ARPES measurements performed on the device shown in Fig. 2, using a photon energy of eV and a beam spot focused to µm2. Figure 3b shows representative angle-integrated and momentum-resolved spectra of this flake. In particular, the dispersion of the main valence bands, formed by S orbitals 34, is well resolved (see also Fig. S3). This is further validated by a constant energy contour at eV with details and resolution comparable to those in studies on bulk crystals 58, 59, 60, 40, 61.
@@
A spatially resolved Ta core-level scan with a step size of µm is used to group the data points based on spectral similarity by the -means clustering algorithm 62 (see Fig. 3d). This unsupervised learning algorithm partitions the dataset into clusters by minimizing the intra-cluster variance. We determine the optimal using the “elbow” method, where diminishing returns in reducing the within-cluster sum of squares are observed. We obtain (see Fig. S4), where the three clusters correspond to distinct regions of the scanned area: bulk ( layers) and monolayer 1T–TaS2 flake, as well as Au surface, agreeing well with the optical microscope image (see Fig. 3a and 3d, as well as the overlaid image in Fig. S5). All angle-integrated spectra within a cluster are averaged to obtain the mean and standard deviation shown in Fig. 3e. A Shirley-type background 63 is subtracted to account for inelastic scattering (see Fig. S6).
@@
Within the bulk-like cluster, the Ta core levels exhibit a splitting stemming from the different electronic environments of Ta atoms within the star-of-David-shaped lattice distortion in the CDW phase 33. As shown in the inset in Fig. 3e, there are three distinct Ta sites, namely a central (A), as well as an inner (B) and outer (C) ring position. Due to their different electronic environment and the charge redistribution towards the central A site, the Ta core levels appear at different binding energies, i.e. eV (C), eV (B), and eV (A) in accordance with measurements of bulk single-crystals 64. The ratio of the integrated intensities matches that of the Ta atoms of each species, i.e. ABC 166. We attribute deviations of the fit from the experimental data, most pronounced for the C species, to inequivalent C sites due to the complex out-of-plane stacking 31, 32, or puckering distortions 65 present in the CDW states of 1T–TaS2. The discrepancy at the edge of the spectrum can be attributed to systematic errors from the background subtraction.
@@
In contrast, the spectrum taken from the monolayer cluster shows a reduction of the splitting between the B and C species, alongside a reduced relative C intensity, which we attribute to the absence of out-of-plane stacking. A shift of eV towards lower binding energies is observed, which we ascribe to charge transfer from Au to 1T–TaS2 and is consistent with findings in monolayer MoS2 on Au contacts 24. To study the influence of 1T–TaS2 / Au interface in more detail, one could aim for a device with a 1T–TaS2 monolayer that extends over the gapped region of the substrate in a future experiment, allowing for a direct comparison of the Au-bound and freestanding monolayer.
3 Conclusions